Experts claim that 10% of murders are committed by women. Is there enough evidence to reject the claim if in a sample of 67 murders, 10 were committed by women? At , test the hypothesis using the p-value method



1
Expert's answer
2021-10-13T11:21:29-0400

"H_0:p=0.1."

"H_a:p\ne0.1."

Test statistic: "z=\frac{\frac{10}{67}-0.1}{\sqrt{\frac{0.1(1-0.1)}{67}}}=1.34."

P-value: "p=2P(Z>1.34)=0.1802."

Since the p-value is greater than 0.05, fail to reject the null hypothesis.

There is not enough evidence to reject the claim that 10% of murders are committed by women.